### v3

# A blow-up result for a wave equation with critical Glassey derivative-type nonlinearity and not Dini continuous modulus of continuity

Abstract
In the present manuscript, we consider a semilinear wave equation with derivative-type nonlinearity. More precisely, we consider a nonlinear term depending on the time derivative of the solution, which is the product of a power nonlinearity with critical Glassey exponent and a not Dini continuous modulus of continuity. By employing Zhou’s approach along a certain characteristic line, we prove the blow-up in finite time for classical solutions (under suitable sign condition for the Cauchy data) and we derive upper bound estimates for the lifespan.
Keywords: semilinear wave equation, Glassey exponent, blow-up, modulus of continuity, Dini condition, lifespan estimates
AMS Classification (2020) 35B44, 35L05, 35L71
1 Introduction
Over the past forty years, several authors studied the Cauchy problem for semilinear wave equations with derivative-type nonlinearity
$\begin{cases}u_{tt}-\Delta u=|u_{t}|^{p},&x\in\mathbb{R}^{n},\ t>0,\\ u(0,x)=\varepsilon u_{0}(x),\ u_{t}(0,x)=\varepsilon u_{1}(x),&x\in\mathbb{R}^{n},\end{cases}$ (1.1)
where the exponent of the semilinear term satisfies $p>1$ and $\varepsilon>0$ is a parameter describing the size of the Cauchy data. Nowadays, it is know that the critical exponent for the semilinear Cauchy problem (1.1) is given by the so-called Glassey exponent
$p_{\mathrm{Gla}}(n):=\frac{n+1}{n-1}$ (1.2)
for $n\geqslant 2$. By critical exponent here we mean a threshold value in the range for the power of the nonlinear term $|u_{t}|^{p}$ such that for $1<p<p_{\mathrm{Gla}}(n)$ (sub-critical case) local in time solutions blow up in finite time under suitable sign assumptions for the Cauchy data and regardless of their size, while for $p>p_{\mathrm{Gla}}(n)$ (super-critical case) solutions are globally in time defined (in a suitable function space) provided that the Cauchy data are sufficiently small. For this particular model, the critical case $p=p_{\mathrm{Gla}}(n)$ belongs to the blow-up range as well. The exponent $p_{\mathrm{Gla}}(n)$ is named after Robert T. Glassey, who conjectured the exact value of the critical exponent (in the sense we just explained) for (1.1) in [12], a Mathematical Review to the paper [36]. Furthermore, since in the 1-dimensional case local solutions to (1.1) blow up for any $p>1$ (under sign assumptions for the data), it is custom to set $p_{\mathrm{Gla}}(1):=+\infty$. Moreover, it is known that for $1<p\leqslant p_{\mathrm{Gla}}(n)$ the lifespan $T_{\varepsilon}$ of a local solution (i.e. the maximal existence time for the solution) satisfies the following sharp estimates:
$T_{\varepsilon}\simeq\begin{cases}C\varepsilon^{-\left(\frac{1}{p-1}-\frac{n-1}{2}\right)^{-1}}&\mbox{if}\ 1<p<p_{\mathrm{Gla}}(n),\\ \exp\big{(}C\varepsilon^{-(p-1)}\big{)}&\mbox{if}\ p=p_{\mathrm{Gla}}(n),\end{cases}$ (1.3)
provided that $\varepsilon\in(0,\varepsilon_{0}]$ for some $\varepsilon_{0}=\varepsilon_{0}(n,p,u_{0},u_{1})>0$.
Let us outline briefly the main contributions in the proof of the above mentioned Glassey’s conjecture and in the derivation of the sharp lifespan estimates in the sub-critical and critical cases (we are going to emphasize in each case whether the lifespan estimates are obtained from above or from below). The study of (1.1) was initiated by Fritz John, who proved that for $p=2$ classical solutions to (1.1) blow-up in the 3-dimensional case in [21], among other results for more general nonlinear terms. Then, in [36] it was proved the global existence of small data and radial solutions in the super-critical case $p>2$ for $n=3$. In [29], John’s result for $p=2$ is extended to space dimensions $n=1,2$. The blow-up results in space dimension $n=2$ were proved for the first time in [35, 1] for the critical case ($p=3$) and the sub-critical case ($31<p<$ 3), respectively. For higher space dimensions ($n\geqslant 4$) blow-up results were proved for radially symmetric solutions in [34]. The global existence in the super-critical case for $n=2,3$ without radial symmetry was proved independently in [18, 39]. Then, Yi Zhou [42] introduced a simpler proof of the blow-up results for classical solutions and for any $n\geqslant 1$; moreover, he derived sharp upper bound estimates for the lifespan in the general not radially symmetric case. The global existence in the super-critical case for $n\geqslant 4$ is proved in [19] for radially symmetric solutions. Furthermore, in [19] lower bound estimates for the lifespan are derived in the sub-critical ($n\geqslant 2$) and critical case ($n\geqslant 3$). The lower bound estimates for the lifespan in the critical case for $n=2$ are derived in [10]. The upper bound estimates and the blow-up results for weak solutions are obtained with a modified test function method in [20] for all space dimensions. Finally, quite recently, the lower bound estimates were proved in the 1-dimensional case (as special unweighted case) in [22].
In the last decade, the problem considered in Glassey’s conjecture has been also investigated for some hyperbolic equations that are different from the classical wave equation on the Minkowski spacetime. On the one hand, we have results for wave equations on manifolds equipped with non-flat metrics: on asymptotically Euclidean manifolds [40, 26] and on Friedmann-Lemaître-Robertson-Walker spacetimes [38, 15, 16]. On the other hand, several Glassey-type exponents have been established from the necessary part in semilinear wave models with time-dependent coefficients: we recall the results for damped wave equations with summable time-dependent coefficients for the linear damping terms [25, 31], for the wave equation with scale-invariant damping and mass terms [33, 13] and for the Tricomi equation [28, 24, 14].
We can rephrase the critical nature of the Glassey exponent for (1.1) by saying that $p=p_{\mathrm{Gla}}(n)$ is the threshold value that separates the blow-up range from the small data solutions’ global existence range in the scale of power nonlinearities $\{|u_{t}|^{p}\}_{p>1}$. Following the ideas of [8], we wonder what happens if we consider the scale of nonlinear terms $|u_{t}|^{p_{\mathrm{Gla}}(n)}\mu(|u_{t}|)$, where $\mu:[0,+\infty)\to[0,+\infty)$ is a modulus of continuity, i.e., $\mu$ is a continuous, concave and increasing function such that $\mu(0)=0$. In other words, we consider the semilinear Cauchy problem for the wave equation with derivative-type nonlinearity
$\begin{cases}u_{tt}-\Delta u=|u_{t}|^{p_{\mathrm{Gla}}(n)}\mu(|u_{t}|),&x\in\mathbb{R}^{n},\ t>0,\\ u(0,x)=\varepsilon u_{0}(x),\ u_{t}(0,x)=\varepsilon u_{1}(x),&x\in\mathbb{R}^{n},\end{cases}$ (1.4)
for $n\geqslant 2$. The additional factor $\mu(|u_{t}|)$ on the right-hand side of the equation in (1.4) provides an additional regularity to the nonlinear term. In particular, for Hölder’s moduli of continuity, i.e., for the class of moduli of continuity $\{\tau^{\beta}\}_{\beta\in(0,1]}$, the previous considerations tell us that any Hölder’s modulus of continuity belongs to the class of moduli of continuity that guarantees the global existence of small data solutions. Our goal in the present manuscript is to prescribe a condition on $\mu$ so that the classical local solutions to (1.4) blow up in finite (under certain assumptions for the Cauchy data). In this sense, our aim is to show that the scale $\{|u_{t}|^{p}\}_{p>1}$ is too rough to describe the critical regularity for the nonlinear term $|u_{t}|^{p_{\mathrm{Gla}}(n)}\mu(|u_{t}|)$. Indeed, we are going to prove that for a modulus of continuity that does not satisfy the Dini condition (cf. Definition 2.1), classical local in time solutions cannot be globally in time prolonged if the moment of order 0 of the second data is positive (see Theorem 2.1). Clearly, since we need to consider $\mu$ which gives a regularity weaker than any Hölder’s regularity, we shall consider a $Log^{-\alpha}$-regularity with $\alpha>0$, or an even weaker regularity. For the nomenclature of the moduli of continuity we address the reader to the introduction of [27].
Next, we shortly review the results for semilinear wave equations having as nonlinear term a critical power nonlinearity multiplied by an additional factor involving a modulus of continuity weaker than any Hölder’s modulus of continuity. In [8], the critical regularity for the semilinear damped wave equation is investigated. To the knowledge of the authors, [8] is the first paper where the problem of determining the critical nonlinearity is studied for a Cauchy problem associated to a wave model. More precisely, the nonlinear term $|u|^{p_{\mathrm{Fuj}}(n)}\mu(|u|)$ is considered in [8] and a sharp integral condition for $\mu$ in a neighborhood of 0 is derived (for $n=1,2$ in the sufficiency part). We stress that for the semilinear damped wave equation the critical exponent for the power nonlinearity is the so-called Fujita exponent $p_{\mathrm{Fuj}}(n)=1+2/n$ and coincides with the critical exponent for the corresponding Cauchy problem for the semilinear heat equation. Remarkably, the integral condition on $\mu$ which provides the blow-up result in [8] coincides with the one that we prescribe in our main result.
Later, in [6] these results are generalized to the corresponding weakly coupled system. In [30] the problem of determining the critical nonlinearity is investigated in the 1-dimensional case for semilinear wave equations either with an effective or a scale-invariant damping term. In some sense, it is not surprising that the threshold condition for $\mu$ is the same one as in [8] in both cases, since the damping structure in these cases is similar to the one for the classical damping (the critical exponent is $p_{\mathrm{Fuj}}(1)=3$). Recently, several semilinear evolution models are considered in [11] with the nonlinear terms given by the product of a critical power nonlinearity and a modulus of continuity, provided that the critical exponent for the power nonlinearity is obtained by the scaling properties of the partial differential operator. We underline that for the semilinear models studied in [11] the presence of linear damping terms plays a crucial role in the asymptotic behaviors of the solutions to the corresponding linearized problems.
In [5] the problem for the critical regularity $|u|^{p_{\mathrm{Str}}(n)}\mu(|u|)$ for the semilinear classical wave equations is considered. Here, $p_{\mathrm{Str}}(n)$ stands for the Strauss exponent (see, for example, the introduction of [41] and references therein for a historical overview on the Strauss’ conjecture, named after the author of [37]). The exponent $p_{\mathrm{Str}}(n)$ is the positive root of the quadratic equation $(n-1)p^{2}-(n+1)p-2=0$ (as for the Glassey exponent, in the 1-dimensional case we set $p_{\mathrm{Str}}(1)=+\infty$) and is the critical exponent for the Cauchy problem associated to the semilinear wave equation $u_{tt}-\Delta u=|u|^{p}$. In particular, by working with an additional $Log^{-\gamma}$ regularity through the modulus of continuity in the nonlinear term $|u|^{p_{\mathrm{Str}}(n)}\mu(|u|)$, in [5] it is proved that in the sub-critical case $\gamma\in(0,1/p_{\mathrm{Str}}(n))$ energy solutions blow up in finite time (under certain assumptions for the data) and that in the 3-dimensional and radially symmetric case this threshold is sharp, since for $\gamma>1/p_{\mathrm{Str}}(3)$ the global existence of small data solutions holds in weighted $L^{\infty}$ spaces.
In this paper, we are going to determine an integral condition on the modulus of continuity $\mu$ so that a blow-up result holds true for (1.4). We are going to adapt Zhou’s approach from [42] to the model in (1.4). This method consists in proving the blow-up on a characteristic line for a classical solution and it has been applied to several semilinear hyperbolic models with derivative-type nonlinearity, cf. [33, 28, 15, 32]. As byproduct of this technique, we also obtain upper bound estimates for the lifespan for small data.
Notation
We write $f\lesssim g$ if there exists a positive constant $C$ such that $f\leqslant Cg$. The relation $f\simeq g$ holds if and only if $g\lesssim f\lesssim g$. When writing $h(\tau)>0$ [resp. $h(\tau)<0$] as $\tau\to 0^{+}$, we mean that there exists $\tau_{0}>0$ such that $h(\tau)>0$ [resp. $h(\tau)<0$] for any $\tau\in(0,\tau_{0}]$. Moreover, $B_{R}$ denotes the ball around the origin with radius $R$ in the whole space and $\omega_{n}:=\frac{1}{\Gamma(\frac{n}{2}+1)}\pi^{\frac{n}{2}}$ is the $n$-dimensional measure of the unit ball $B_{1}$ with Euler’s gamma function $\Gamma$. The average of $f$ over $\Omega$ with respect to a measure $\sigma$ on $\Omega$ is denoted by
[…]
Finally, $[y]_{+}:=\max\{y,0\}$ is the positive part of $y\in\mathbb{R}$ and for any $k\in\mathbb{N}^{*}$ we denote by $\log^{[k]}$ and $\exp^{[k]}$ the $k$-times iterated logarithmic and exponential functions, respectively, that is
$\log^{[k]}:=\begin{cases}\log&\mbox{if}\ k=1,\\ \log\circ\log^{[k-1]}&\mbox{if}\ k\geqslant 2,\end{cases}$ $\exp^{[k]}:=\begin{cases}\exp&\mbox{if}\ k=1,\\ \exp\circ\exp^{[k-1]}&\mbox{if}\ k\geqslant 2.\end{cases}$
2 Main results
Let us begin by recalling the notion of Dini condition (or Dini test) for a modulus of continuity.
Definition 2.1.
Let $\mu:[0,+\infty)\to[0,+\infty)$ be a modulus of continuity, that is, $\mu$ is a continuous, increasing, concave function and satisfies $\mu(0)=0$. Then, $\mu$ satisfies the Dini condition if
$\int_{0}^{\tau_{0}}\frac{\mu(\tau)}{\tau}\mathrm{d}\tau<+\infty\qquad\mbox{for some}\ \tau_{0}>0.$ (2.1)
Remark 2.1.
Notice that if (2.1) is satisfied for some $\tau_{0}>0$ then it holds for any $\tau_{0}>0$. Indeed, being $\mu$ continuous, increasing and $\mu(0)=0$, the function $\frac{\mu(\tau)}{\tau}$ is locally summable in any compact set $K\subset(0,+\infty)$.
The condition in (2.1) was implicitly introduced in the study of the convergence of the Fourier series in [7, Page 101] (see also [43]). We emphasize that, albeit different definitions are possible for the notion of modulus of continuity, the Dini condition appears in different areas of modern mathematical analysis: from elliptic equations [17, 3, 23] and integral equations [4], to dynamical systems [9]. We refer to [2, Section 2, Paragraph A.2] for a class of non-trivial examples of Dini-moduli of continuity obtained through generalized power series.
After introducing the Dini condition, we can state our blow-up result for semilinear classical wave equations with derivative type critical nonlinearity.
Theorem 2.1.
Let $n\geqslant 2$ and let $u_{0}\in\mathcal{C}^{2}_{0}(\mathbb{R}^{n})$, $u_{1}\in\mathcal{C}^{1}_{0}(\mathbb{R}^{n})$ be compactly supported functions with supports contained in $B_{R}$ for some $R>0$. Furthermore, we assume that $u_{1}$ satisfies
$\int_{\mathbb{R}^{n}}u_{1}(x)\,\mathrm{d}x>0.$ (2.2)
Let $\mu:[0,+\infty)\to[0,+\infty)$ be a modulus of continuity fulfilling the following assumptions:
  • $(A_{1})$
    the function $g:\tau\in\mathbb{R}\to|\tau|^{p_{\mathrm{Gla}}(n)}\mu(|\tau|)\in[0,+\infty)$ is convex;
  • $(A_{2})$
    $\mu$ does not satisfies the Dini condition, i.e., for some (and, then, for any) $\tau_{0}>0:$ $\frac{\mu(\tau)}{\tau}\not\in L^{1}([0,\tau_{0}])$, namely,
    $\int_{0}^{\tau_{0}}\frac{\mu(\tau)}{\tau}\,\mathrm{d}\tau=+\infty.$ (2.3)
If $u\in\mathcal{C}^{2}([0,T_{\varepsilon})\times\mathbb{R}^{n})$ is a classical local (in time) solution to the semilinear Cauchy problem (1.4), then $u$ blows up in finite time, i.e., $T_{\varepsilon}<+\infty$.
Remark 2.2.
We may rephrase the assumption on the nonlinear term $|u_{t}|^{p_{\mathrm{Gla}}(n)}\mu(|u_{t}|)$ in Theorem 2.1 by saying that $\mu$ is a modulus of continuity not Dini continuous.
Example 2.1.
Let $\tau_{0}>0$ be sufficiently small. The assumptions $(A_{1})-(A_{2})$ on the modulus of continuity (MoC) from Theorem 2.1 hold for the following functions $\mu(\tau)$ that on a small interval $[0,\tau_{0}]$ satisfy $\mu(0)=0$ and for $\tau\in(0,\tau_{0}]$:
$\mu(\tau)=(\log\frac{1}{\tau})^{-\gamma}$ with $\gamma\in(0,1]$;
${\mu(\tau)=\prod_{j=1}^{k-1}\left(\log^{[j]}\tfrac{1}{\tau}\right)^{-1}\left(\log^{[k]}\tfrac{1}{\tau}\right)^{-\gamma}}$ with $\gamma\in(0,1]$ and $k\in\mathbb{N}$, $k\geqslant 2$;
${\mu(\tau)=\prod_{j=1}^{k}\left(\log^{[j]}\tfrac{1}{\tau}\right)^{-\gamma_{j}}}$ with $k\in\mathbb{N}$, $k\geqslant 2$ and $\{\gamma_{j}\}_{1\leqslant j\leqslant k}\subset\mathbb{R}$ fulfilling the next conditions ((((clearly, we are excluding the trivial case $\gamma_{j}=0$ for any $j\in\{1,\ldots,k\})$
  1. a))))
    $\gamma_{j_{0}}>0$ where $j_{0}:=\min\{1\leqslant j\leqslant k:\gamma_{j}\neq 0\}$;
  2. b))))
    either $\gamma_{j}=1$ for any $j\in\{1,\ldots,k\}$ or $\gamma_{j_{1}}<1$, where $j_{1}:=\min\{1\leqslant j\leqslant k:\gamma_{j}\neq 1\}$ ((((of course, in this second occurrence $\gamma_{j}\neq 1$ for some $j\in\{1,\ldots,k\})$.
Remark 2.3.
We consider the most general case in Example 2.1, i.e., $(3\mathrm{MoC})$, and we explain why $\mu$ is a modulus of continuity that satisfies $(A_{1})-(A_{2})$. The continuity in $\tau=0$ follows from the fact that $\gamma_{j_{0}}>0$. For the monotonicity and the concavity of $\mu$ in a right neighborhood of 00 we need to calculate $\mu^{\prime}(\tau)$ and $\mu^{\prime\prime}(\tau)$. Since
$\frac{\mathrm{d}}{\mathrm{d}\tau}\left(\log^{[h]}\frac{1}{\tau}\right)=-\frac{1}{\tau}\prod_{j=1}^{h-1}\left(\log^{[j]}\frac{1}{\tau}\right)^{-1},$
by Leibniz rule we find
$\mu^{\prime}(\tau)$ $=\frac{\mu(\tau)}{\tau}\sum_{h=1}^{k}\gamma_{h}\prod_{j=1}^{h}\left(\log^{[j]}\frac{1}{\tau}\right)^{-1},$ (2.4)
$\mu^{\prime\prime}(\tau)$ $=\frac{\mu(\tau)}{\tau^{2}}\bigg{[}-\sum_{h=1}^{k}\gamma_{h}\prod_{j=1}^{h}\left(\log^{[j]}\frac{1}{\tau}\right)^{-1}+\sum_{h=1}^{k}\gamma_{h}\sum_{\ell=1}^{h}(\gamma_{\ell}+1)\prod_{j=1}^{\ell}\left(\log^{[j]}\frac{1}{\tau}\right)^{-2}\prod_{j=\ell+1}^{h}\left(\log^{[j]}\frac{1}{\tau}\right)^{-1}$
$\qquad\qquad+\sum_{h=1}^{k}\gamma_{h}\prod_{j=1}^{h}\left(\log^{[j]}\frac{1}{\tau}\right)^{-2}\sum_{\ell=h+1}^{k}\gamma_{\ell}\prod_{j=h+1}^{\ell}\left(\log^{[j]}\frac{1}{\tau}\right)^{-1}\bigg{]}.$ (2.5)
Since $\gamma_{j_{0}}>0$ it follows that $\mu^{\prime}(\tau)>0$ and $\mu^{\prime\prime}(\tau)<0$ as $\tau\to 0^{+}$. Hence, $\mu$ is strictly increasing and concave in $[0,\tau_{0}]$, provided that $\tau_{0}$ is sufficiently small. We check now that the moduli of continuity in Example 2.1 fulfill $(A_{1})$. By (2.4) and (2.5) we see that
$\tau^{k}\mu^{(k)}(\tau)=o(\mu(\tau))\qquad\mbox{as}\ \tau\to 0^{+}\ \mbox{for}\ k=1,2.$ (2.6)
Consequently, recalling that $g(\tau)=\tau^{p_{\mathrm{Gla}}(n)}\mu(\tau)$ for $\tau>0$, we obtain
$g^{\prime\prime}(\tau)$ $=\tau^{p_{\mathrm{Gla}}(n)-2}\left[p_{\mathrm{Gla}}(n)(p_{\mathrm{Gla}}(n)-1)\mu(\tau)+2p_{\mathrm{Gla}}(n)\tau\mu^{\prime}(\tau)+\tau^{2}\mu^{\prime\prime}(\tau)\right]$
$=p_{\mathrm{Gla}}(n)(p_{\mathrm{Gla}}(n)-1)\tau^{p_{\mathrm{Gla}}(n)-2}\mu(\tau)\left[1+o(1)\right].$
Thus, $g^{\prime\prime}(\tau)>0$ as $\tau\to 0^{+}$ and $g$ is convex in $[0,\tau_{0}]$, provided that $\tau_{0}$ is sufficiently small. Let us show that the moduli of continuity in Example 2.1 satisfy $(A_{2})$. If $\gamma_{j}=1$ for any $j\in\{1,\ldots,k\}$, obviously
$\int_{0}^{\tau_{0}}\frac{\mu(\tau)}{\tau}\mathrm{d}\tau=\int_{0}^{\tau_{0}}\tau^{-1}\prod_{j=1}^{k}\left(\log^{[j]}\frac{1}{\tau}\right)^{-1}\mathrm{d}\tau=\lim_{\epsilon\to 0^{+}}\left(\log^{[k+1]}\frac{1}{\epsilon}-\log^{[k+1]}\frac{1}{\tau_{0}}\right)=+\infty$
On the other hand, if $\gamma_{j_{1}}<1$, then we can fix $\delta>0$ such that $\gamma_{j_{1}}+\delta<1$. Since
$\lim_{\tau\to 0^{+}}\left(\log^{[j_{1}]}\frac{1}{\tau}\right)^{\delta}\prod_{j=j_{1}+1}^{k}\left(\log^{[j]}\frac{1}{\tau}\right)^{-\gamma_{j}}=+\infty,$
being $\tau_{0}$ sufficiently small, it results
$\int_{0}^{\tau_{0}}\frac{\mu(\tau)}{\tau}\mathrm{d}\tau$ $\gtrsim\int_{0}^{\tau_{0}}\tau^{-1}\prod_{j=1}^{j_{1}-1}\left(\log^{[j]}\frac{1}{\tau}\right)^{-1}\left(\log^{[j_{1}]}\frac{1}{\tau}\right)^{-\gamma_{j_{1}}-\delta}\mathrm{d}\tau$
$=\frac{1}{1-\gamma_{j_{1}}-\delta}\lim_{\epsilon\to 0^{+}}\left(\left(\log^{[j_{1}]}\frac{1}{\epsilon}\right)^{1-\gamma_{j_{1}}-\delta}-\left(\log^{[j_{1}]}\frac{1}{\tau_{0}}\right)^{1-\gamma_{j_{1}}-\delta}\right)=+\infty.$
Finally, we point out that for a modulus of continuity $\mu$ it is crucial the behavior on a right neighborhood of 00 ((((for this reason in the previous considerations we focused on the interval $[0,\tau_{0}])$ and that $\mu(\tau)$ can be continued to $\tau\in[0,+\infty)$ in such a way that $\mu$ is a continuous, concave and increasing function and the function $g$ ((((that provides the nonlinear term)))) is a convex function. Moreover, since we prolong $g$ to an increasing and convex function in $[0,+\infty)$, by even reflection we obtain that $g$ is convex in $\mathbb{R}$. As we will see in the proof of Theorem 2.1, the behavior of $\mu(\tau)$ [[[[resp. $g(\tau)]$ for large $\tau$ is completely irrelevant while proving the blow-up of a local solution to (1.4).
Example 2.2.
As a counterexample for the condition (2.3) in Theorem 2.1 we consider the family of Hölder’s moduli of continuity, that is, $\{\mu_{\beta}\}_{\beta\in(0,1]}$ such that $\mu_{\beta}(\tau)=\tau^{\beta}$ for any $\tau\in[0,+\infty)$ and $\beta\in(0,1]$. Therefore, this condition is consistent with the global (in time) existence results in [36, 18, 39, 19] for the classical semilinear wave equations with derivative-type nonlinearity $|u_{t}|^{p_{\mathrm{Gla}}(n)+\beta}$, which correspond to the super-critical case $p>p_{\mathrm{Gla}}(n)$ in (1.1).
Remark 2.4.
Concerning the semilinear wave equation (1.1) in the critical case $p=p_{\mathrm{Gla}}(n)$, the previous works [21, 29, 35, 34, 1, 42] showed that the solution blows up in finite time. By taking $\mu(|u_{t}|)$ as an additional factor that multiplies the critical power nonlinearity $|u_{t}|^{p_{\mathrm{Gla}}(n)}$, for a modulus of continuity $\mu$ fulfilling (2.3), Theorem 2.1 tells us that a classical solution still blows up in finite time. Notice that we have to consider moduli of continuity weaker than any Hölder’s modulus of continuity, therefore, we look for an infinitesimal function $\mu(\tau)$ as $\tau\to 0^{+}$ but that tends to 00 slower than any power $\{\tau^{\beta}\}_{\beta\in(0,1]}$.
As byproduct of our approach in proving the blow-up result, we derive an upper bound estimate for the lifespan $T_{\varepsilon}$ of a local solution considered in Theorem 2.1.
Proposition 2.1.
Let $n\geqslant 2$. We require that the initial data $u_{0},u_{1}$ and the modulus of continuity $\mu$ satisfy the same assumptions as in Theorem 2.1. Let us introduce the decreasing function
$\mathcal{H}:\tau\in(0,\tau_{0}]\longrightarrow\int_{\tau}^{\tau_{0}}\frac{\mu(\sigma)}{\sigma}\mathrm{d}\sigma\in[0,+\infty).$
Then, there exists $\varepsilon_{0}=\varepsilon_{0}(n,R,u_{1})>0$ such that for any $\varepsilon\in(0,\varepsilon_{0}]$ the following upper bound estimate for the lifespan of the classical local (in time) solution $u$ holds:
$T_{\varepsilon}\leqslant C\varepsilon^{\frac{2}{n-1}}\left[\mathcal{H}^{-1}\left(c_{1}\,\varepsilon^{-\frac{2}{n-1}}+\mathcal{H}(c_{2}\varepsilon)\right)\right]^{-\frac{2}{n-1}},$ (2.7)
where the positive constants $c_{1},c_{2},C$ are independent of $\varepsilon$.
In Section 4, we rewrite (2.7) in the special cases $(1\mathrm{MoC})$ and $(2\mathrm{MoC})$ from Example 2.1.
3 Blow-up of the classical solutions
To begin with the proof, let us introduce the notation $x=(z,w)\in\mathbb{R}^{n}$ with $z\in\mathbb{R}$ and $w\in\mathbb{R}^{n-1}$ (we recall that we are considering the case $n\geqslant 2$). Then, we may define the next auxiliary functions:
$\mathcal{U}(t,z)$ $:=\int_{\mathbb{R}^{n-1}}u(t,z,w)\,\mathrm{d}w\qquad\mbox{for any}\ (t,z)\in[0,T_{\varepsilon})\times\mathbb{R},$
$\mathcal{U}_{j}(z)$ $:=\int_{\mathbb{R}^{n-1}}u_{j}(z,w)\,\mathrm{d}w\qquad\ \mbox{forany}\ z\in\mathbb{R}\ \mbox{and}\ j=0,1.$
The function $\mathcal{U}=\mathcal{U}(t,z)$ solves the following nonlinear wave equations in 1d:
$\begin{cases}\displaystyle{\mathcal{U}_{tt}-\mathcal{U}_{zz}=\int_{\mathbb{R}^{n-1}}|u_{t}(t,z,w)|^{p_{\mathrm{Gla}}(n)}\mu(|u_{t}(t,z,w)|)\,\mathrm{d}w,}&z\in\mathbb{R},\ t\in(0,T_{\varepsilon}),\\ \mathcal{U}(0,z)=\varepsilon\mathcal{U}_{0}(z),\ \mathcal{U}_{t}(0,z)=\varepsilon\mathcal{U}_{1}(z),&z\in\mathbb{R}.\end{cases}$
By D’Alembert’s formula, recalling $g(\tau)=|\tau|^{p_{\mathrm{Gla}}(n)}\mu(|\tau|)$, we obtain the representation
$\mathcal{U}(t,z)=\varepsilon\mathcal{U}^{\operatorname{lin}}(t,z)+\mathcal{U}^{\operatorname{nlin}}(t,z),$ (3.1)
where
$\mathcal{U}^{\operatorname{lin}}(t,z)$ $:=\frac{1}{2}\big{(}\mathcal{U}_{0}(z+t)+\mathcal{U}_{0}(z-t)\big{)}+\frac{1}{2}\int_{z-t}^{z+t}\mathcal{U}_{1}(y)\,\mathrm{d}y,$
$\mathcal{U}^{\operatorname{nlin}}(t,z)$ $:=\frac{1}{2}\int_{0}^{t}\int_{z-t+s}^{z+t-s}\int_{\mathbb{R}^{n-1}}g\big{(}u_{t}(s,y,w)\big{)}\,\mathrm{d}w\,\mathrm{d}y\,\mathrm{d}s.$ (3.2)
3.1 Nonlinear integral inequality with the modulus of continuity
Our next goal is to determine a nonlinear integral inequality for the function $\mathcal{U}(t,z)$ on the characteristic line $t-z=R$ involving the modulus of continuity. Due to the property of finite speed of propagation, the assumption $\mathrm{supp}\,u_{j}\subset B_{R}$ for $j=0,1$ implies that
$\mathrm{supp}\,u(t,\cdot)\subset B_{R+t}\qquad\mbox{for any}\ t\in[0,T_{\varepsilon}).$
Consequently, for $z\in\mathbb{R}$ such that $|z|\geqslant R+t$ we have that the integrand function in the definition of $\mathcal{U}(t,z)$ satisfies $u(t,z,w)=0$ for any $w\in\mathbb{R}^{n-1}$, because $|(z,w)|^{2}\geqslant(R+t)^{2}+|w|^{2}\geqslant(R+t)^{2}$. Hence,
$\mathrm{supp}\,\mathcal{U}(t,\cdot)\subset\big{(}-(R+t),R+t\big{)}\ \qquad\mbox{for any}\ t\in[0,T_{\varepsilon}).$ (3.3)
Since our nonlinear term $g(u_{t})$ is local, if we rewrite the support condition with respect to the variable $w\in\mathbb{R}^{n-1}$, we find
$\mathrm{supp}\,g\big{(}u_{t}(s,y,\cdot)\big{)}\subset\mathrm{supp}\,u_{t}(s,y,\cdot)\subset\left\{w\in\mathbb{R}^{n-1}:\ |w|^{2}\leqslant(R+s)^{2}-y^{2}\right\}$
for any $s\in[0,t]$ and any $y\in\mathbb{R}$ such that $|y|\leqslant R+s$. Consequently, for any $y\in\mathbb{R}$ such that $|y|\leqslant R+s$ we have
$\int_{\mathbb{R}^{n-1}}g\big{(}u_{t}(s,y,w)\big{)}\,\mathrm{d}w$ $=\int_{B_{\sqrt{(R+s)^{2}-y^{2}}}}g\big{(}u_{t}(s,y,w)\big{)}\,\mathrm{d}w$
$=\omega_{n-1}\big{(}(R+s)^{2}-y^{2}\big{)}^{\frac{n-1}{2}}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B_{\sqrt{(R+s)^{2}-y^{2}}}}g\big{(}u_{t}(s,y,w)\big{)}\,\mathrm{d}w.$
Thanks to the assumption $(A_{1})$, which guarantees the convexity $g$, by Jensen’s inequality we get
$\int_{\mathbb{R}^{n-1}}g\big{(}u_{t}(s,y,w)\big{)}\,\mathrm{d}w$ $\geqslant\omega_{n-1}\big{(}(R+s)^{2}-y^{2}\big{)}^{\frac{n-1}{2}}g\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B_{\sqrt{(R+s)^{2}-y^{2}}}}u_{t}(s,y,w)\,\mathrm{d}w\right)$
$=\omega_{n-1}\big{(}(R+s)^{2}-y^{2}\big{)}^{\frac{n-1}{2}}g\left(\frac{1}{\omega_{n-1}}\big{(}(R+s)^{2}-y^{2}\big{)}^{-\frac{n-1}{2}}\mathcal{U}_{t}(s,y)\right)$
for any $s\in[0,t]$ and any $y\in[z-t+s,z+t-s]$ such that $|y|\leqslant R+s$. From (3.2) we may estimate
$\mathcal{U}^{\operatorname{nlin}}(t,z)$ $\geqslant\frac{\omega_{n-1}}{2}\int_{0}^{t}\int_{z-t+s}^{z+t-s}\big{[}(R+s)^{2}-y^{2}\big{]}_{+}^{\frac{n-1}{2}}g\left(\frac{1}{\omega_{n-1}}\big{[}(R+s)^{2}-y^{2}\big{]}_{+}^{-\frac{n-1}{2}}\mathcal{U}_{t}(s,y)\right)\mathrm{d}y\,\mathrm{d}s$
$=\frac{\omega_{n-1}}{2}\int_{z-t}^{z+t}\int_{0}^{t-|z-y|}\big{[}(R+s)^{2}-y^{2}\big{]}_{+}^{\frac{n-1}{2}}g\left(\frac{1}{\omega_{n-1}}\big{[}(R+s)^{2}-y^{2}\big{]}_{+}^{-\frac{n-1}{2}}\mathcal{U}_{t}(s,y)\right)\mathrm{d}s\,\mathrm{d}y,$
where we used Fubini’s theorem.
Hereafter, we work on the characteristic line $t-z=R$ for $z\geqslant R$ and shrink the domain of integration from the triangular region
[…]
to the parallelogrammatic sub-region
$\left\{(s,y)\in\mathbb{R}^{2}:\ y-R\leqslant s\leqslant y+R,\ R\leqslant y\leqslant z\right\}.$
For this reason, denoting by
$M(y;R):=\int_{y-R}^{y+R}\big{(}(R+s)^{2}-y^{2}\big{)}^{\frac{n-1}{2}}\mathrm{d}s$
the measure of the interval $[y-R,y+R]$ with respect to $\big{(}(R+s)^{2}-y^{2}\big{)}^{\frac{n-1}{2}}\mathrm{d}s$ (when $|y|\leqslant R+s$), it results
$\mathcal{U}^{\operatorname{nlin}}(z+R,z)$ $\geqslant\frac{\omega_{n-1}}{2}\int_{R}^{z}\int_{y-R}^{y+R}g\left(\frac{1}{\omega_{n-1}}\big{(}(R+s)^{2}-y^{2}\big{)}^{-\frac{n-1}{2}}\mathcal{U}_{t}(s,y)\right)\big{(}(R+s)^{2}-y^{2}\big{)}^{\frac{n-1}{2}}\mathrm{d}s\,\mathrm{d}y$
$=\frac{\omega_{n-1}}{2}\int_{R}^{z}M(y;R)\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{y-R}^{y+R}g\left(\frac{1}{\omega_{n-1}}\big{(}(R+s)^{2}-y^{2}\big{)}^{-\frac{n-1}{2}}\mathcal{U}_{t}(s,y)\right)\big{(}(R+s)^{2}-y^{2}\big{)}^{\frac{n-1}{2}}\mathrm{d}s\,\mathrm{d}y.$
Let us derive upper and lower bound estimates for $M(y;R)$ for $y\geqslant R$. On the one hand,
$M(y;R)$ $=\int_{y-R}^{y+R}\big{(}(R+s-y)(R+s+y)\big{)}^{\frac{n-1}{2}}\mathrm{d}s$
$\geqslant(2y)^{\frac{n-1}{2}}\int_{y-R}^{y+R}(R+s-y)^{\frac{n-1}{2}}\mathrm{d}s$
$=\frac{2^{n+1}}{n+1}R^{\frac{n+1}{2}}y^{\frac{n-1}{2}},$
on the other hand,
$M(y;R)$ $\leqslant(2R)^{\frac{n-1}{2}}\int_{y-R}^{y+R}(R+s+y)^{\frac{n-1}{2}}\mathrm{d}s$
$=\frac{2^{n+1}}{n+1}R^{\frac{n-1}{2}}\big{(}(R+y)^{\frac{n+1}{2}}-y^{\frac{n+1}{2}}\big{)}$
$=2^{n}R^{\frac{n+1}{2}}(y+\theta_{0}R)^{\frac{n-1}{2}}$
$\leqslant 2^{\frac{3n-1}{2}}R^{\frac{n+1}{2}}y^{\frac{n-1}{2}},$
for some $\theta_{0}\in(0,1)$, where in the second last step we used the mean value theorem. Summarizing the previous two inequalities, we proved that
$M(y;R)\approx R^{\frac{n+1}{2}}y^{\frac{n-1}{2}}\ \qquad\mbox{as}\ y\geqslant R.$
By applying Jensen’s inequality in the $s$-integral with respect to the measure $\big{(}(R+s)^{2}-y^{2}\big{)}^{\frac{n-1}{2}}\mathrm{d}s$, for $z\geqslant R$ we obtain
$\mathcal{U}^{\operatorname{nlin}}(z+R,z)$ $\geqslant\frac{2^{n}\omega_{n-1}}{n+1}R^{\frac{n+1}{2}}\int_{R}^{z}y^{\frac{n-1}{2}}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{y-R}^{y+R}g\left(\frac{1}{\omega_{n-1}}\big{(}(R+s)^{2}-y^{2}\big{)}^{-\frac{n-1}{2}}\mathcal{U}_{t}(s,y)\right)\big{(}(R+s)^{2}-y^{2}\big{)}^{\frac{n-1}{2}}\mathrm{d}s\,\mathrm{d}y$
$\geqslant\frac{2^{n}\omega_{n-1}}{n+1}R^{\frac{n+1}{2}}\int_{R}^{z}y^{\frac{n-1}{2}}g\left(\frac{1}{\omega_{n-1}}\ \mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{y-R}^{y+R}\mathcal{U}_{t}(s,y)\,\mathrm{d}s\right)\mathrm{d}y$
$=\frac{2^{n}\omega_{n-1}}{n+1}R^{\frac{n+1}{2}}\int_{R}^{z}y^{\frac{n-1}{2}}g\left(\frac{1}{\omega_{n-1}M(y;R)}\int_{y-R}^{y+R}\mathcal{U}_{t}(s,y)\,\mathrm{d}s\right)\mathrm{d}y.$
Since $g$ is an increasing function on $[0,+\infty)$, being the product of two nonnegative and increasing functions, taking
$K:=\omega_{n-1}^{-1}2^{-\frac{3n-1}{2}}R^{-\frac{n+1}{2}},$
we conclude that
$\mathcal{U}^{\operatorname{nlin}}(z+R,z)$ $\geqslant\frac{2^{n}\omega_{n-1}}{n+1}R^{\frac{n+1}{2}}\int_{R}^{z}y^{\frac{n-1}{2}}g\left(Ky^{-\frac{n-1}{2}}\int_{y-R}^{y+R}\mathcal{U}_{t}(s,y)\,\mathrm{d}s\right)\mathrm{d}y$
$=\underbrace{\frac{2^{n}\omega_{n-1}}{n+1}R^{\frac{n+1}{2}}}_{=:C_{1}}\int_{R}^{z}y^{\frac{n-1}{2}}g\left(Ky^{-\frac{n-1}{2}}\mathcal{U}(y+R,y)\right)\mathrm{d}y$ (3.4)
for $z\geqslant R$, where we used the fundamental theorem of calculus associated with $\mathcal{U}(y-R,y)=0$ that follows from the support condition (3.3).
Our assumption on the initial data shows that
$\mathcal{U}^{\operatorname{lin}}(z+R,z)$ $=\frac{1}{2}\big{(}\mathcal{U}_{0}(2z+R)+\mathcal{U}_{0}(-R)\big{)}+\frac{1}{2}\int_{-R}^{2z+R}\mathcal{U}_{1}(y)\,\mathrm{d}y=\frac{1}{2}\int_{\mathbb{R}}\mathcal{U}_{1}(y)\,\mathrm{d}y$
$=\frac{1}{2}\int_{\mathbb{R}^{n}}u_{1}(x)\,\mathrm{d}x=:C_{0}>0,$ (3.5)
since $z\geqslant R$ implies that $2z+R\geqslant R$ and $\mathrm{supp}\,\mathcal{U}_{0},\mathrm{supp}\,\mathcal{U}_{1}\subset(-R,R)$ which is a consequence of the assumption $\mathrm{supp}\,u_{0},\mathrm{supp}\,u_{1}\subset B_{R}$ for the Cauchy data.
Finally, combining (3.4) and (3.5) we obtain the following nonlinear integral inequality:
$\mathcal{U}(z+R,z)\geqslant C_{0}\varepsilon+C_{1}\int_{R}^{z}y^{\frac{n-1}{2}}g\left(Ky^{-\frac{n-1}{2}}\mathcal{U}(y+R,y)\right)\mathrm{d}y.$ (3.6)
3.2 Proof of Theorem 2.1
Let us introduce the $z$-dependent functional
$\mathcal{G}(z):=C_{0}\varepsilon+C_{1}\int_{R}^{z}y^{\frac{n-1}{2}}g\left(Ky^{-\frac{n-1}{2}}\mathcal{U}(y+R,y)\right)\mathrm{d}y,$
for $z\geqslant R$. Clearly, $\mathcal{G}(R)=C_{0}\varepsilon$ and
$\mathcal{G}^{\prime}(z)=C_{1}z^{\frac{n-1}{2}}g\left(Kz^{-\frac{n-1}{2}}\mathcal{U}(z+R,z)\right)>0.$ (3.7)
Hence, $\mathcal{G}=\mathcal{G}(z)$ is an increasing function and, therefore,
$\mathcal{G}(z)\geqslant\mathcal{G}(R)=C_{0}\varepsilon.$ (3.8)
We remark that from (3.6) we have $\mathcal{U}(z+R,z)\geqslant\mathcal{G}(z)$. Consequently, using again the monotonicity of $g$ on the nonnegative semi-axis, we find
$\mathcal{G}^{\prime}(z)$ $\geqslant C_{1}z^{\frac{n-1}{2}}g\left(Kz^{-\frac{n-1}{2}}\mathcal{G}(z)\right)$
$\geqslant C_{2}z^{-\frac{n-1}{2}(p_{\mathrm{Gla}}(n)-1)}[\mathcal{G}(z)]^{p_{\mathrm{Gla}}(n)}\mu\left(Kz^{-\frac{n-1}{2}}\mathcal{G}(z)\right)$ (3.9)
for $z\geqslant R$ with $C_{2}:=C_{1}K^{p_{\mathrm{Gla}}(n)}$.
Recalling the value of $p_{\mathrm{Gla}}(n)$, the inequalities (3.8) and (3.9) yield
$\frac{\mathcal{G}^{\prime}(z)}{[\mathcal{G}(z)]^{p_{\mathrm{Gla}}(n)}}\geqslant C_{2}z^{-1}\mu\left(C_{0}K\varepsilon z^{-\frac{n-1}{2}}\right).$
Integrating the last inequality over $[R,z]$, one arrives at
$\frac{1}{p_{\mathrm{Gla}}(n)-1}\left((C_{0}\varepsilon)^{1-p_{\mathrm{Gla}}(n)}-[\mathcal{G}(z)]^{1-p_{\mathrm{Gla}}(n)}\right)$ $=\int_{R}^{z}\frac{\mathcal{G}^{\prime}(y)}{[\mathcal{G}(y)]^{p_{\mathrm{Gla}}(n)}}\,\mathrm{d}y$
$\geqslant C_{2}\int_{R}^{z}y^{-1}\mu\left(C_{0}K\varepsilon y^{-\frac{n-1}{2}}\right)\mathrm{d}y.$
By using the change of variable $\tau=C_{0}K\varepsilon y^{-\frac{n-1}{2}}$, we derive
$\int_{R}^{z}y^{-1}\mu\left(C_{0}K\varepsilon y^{-\frac{n-1}{2}}\right)\mathrm{d}y=\frac{2}{n-1}\int_{C_{0}Kz^{-\frac{n-1}{2}}\varepsilon}^{C_{0}KR^{-\frac{n-1}{2}}\varepsilon}\frac{\mu(\tau)}{\tau}\,\mathrm{d}\tau.$
That is to say,
$(C_{0}\varepsilon)^{-\frac{2}{n-1}}-[\mathcal{G}(z)]^{-\frac{2}{n-1}}\geqslant\frac{4C_{2}}{(n-1)^{2}}\int_{C_{0}Kz^{-\frac{n-1}{2}}\varepsilon}^{C_{0}KR^{-\frac{n-1}{2}}\varepsilon}\frac{\mu(\tau)}{\tau}\,\mathrm{d}\tau.$ (3.10)
Let us assume that $u$ is global (in time) defined, then $\mathcal{U}(z+R,z)$ and, in turn, $\mathcal{G}(z)$ are defined for any $z\geqslant R$. Nevertheless, this produces a contradiction as $z\to+\infty$, due to the facts that the left-hand side of (3.10) is bounded (recall that thanks to (3.7) and (3.8), $\mathcal{G}(z)$ is an increasing function that is bounded from below by a positive constant), while the right-hand side, by the monotone convergence theorem, tends to
$\frac{4C_{2}}{(n-1)^{2}}\int_{0}^{C_{0}KR^{-\frac{n-1}{2}}\varepsilon}\frac{\mu(\tau)}{\tau}\,\mathrm{d}\tau=+\infty$
since the modulus of continuity $\mu(\tau)$ satisfies the condition (2.3). Our proof of Theorem 2.1 is complete.
3.3 Proof of Proposition 2.1
In order to derive our desired lifespan estimate, let us recall the definition of the function $\mathcal{H}$
$\mathcal{H}(\tau)=\int_{\tau}^{\tau_{0}}\frac{\mu(\sigma)}{\sigma}\,\mathrm{d}\sigma.$
From the previous definition we see that $\mathcal{H}$ is decreasing and ${\lim_{\tau\to 0^{+}}\mathcal{H}(\tau)=+\infty}$ thanks to the assumption (2.3). Consequently, $\mathcal{H}:(0,\tau_{0}]\to[0,+\infty)$ is invertible.
From (3.10) we derive the following inequality:
$[\mathcal{G}(z)]^{-\frac{2}{n-1}}\leqslant(C_{0}\varepsilon)^{-\frac{2}{n-1}}-K_{1}\left(\mathcal{H}(K_{2}z^{-\frac{n-1}{2}}\varepsilon)-\mathcal{H}(c_{2}\varepsilon)\right)$ (3.11)
for $z\geqslant R$, where $c_{2}:=C_{0}KR^{-\frac{n-1}{2}}$, $K_{1}:=\frac{4C_{2}}{(n-1)^{2}}$ and $K_{2}=C_{0}K$. Consequently, for any $z\geqslant R$ such that
$(C_{0}\varepsilon)^{-\frac{2}{n-1}}-K_{1}\left(\mathcal{H}(K_{2}z^{-\frac{n-1}{2}}\varepsilon)-\mathcal{H}(c_{2}\varepsilon)\right)<0$ (3.12)
the inequality in (3.11) is violated (we recall that according to (3.8), $\mathcal{G}(z)$ is a positive function) and, therefore, $\mathcal{G}(z)$, and in turn $\mathcal{U}(R+z,z)$, does not exist. We remind that we are working on the characteristic line $t=R+z$ for $z\geqslant R$, so, $z\leqslant t=z+R\leqslant 2z$. Since the left-hand side of (3.12) is decreasing with respect to $z$, we obtain that $\mathcal{U}(R+z,z)$ cannot be defined whenever $t\geqslant 2R$ fulfills
$(C_{0}\varepsilon)^{-\frac{2}{n-1}}-K_{1}\left[\mathcal{H}\left(K_{2}\left(\frac{t}{2}\right)^{-\frac{n-1}{2}}\varepsilon\right)-\mathcal{H}(c_{2}\varepsilon)\right]<0.$ (3.13)
Then, the lifespan $T_{\varepsilon}$ must satisfy
$(C_{0}\varepsilon)^{-\frac{2}{n-1}}-K_{1}\left[\mathcal{H}\left(K_{3}T_{\varepsilon}^{-\frac{n-1}{2}}\varepsilon\right)-\mathcal{H}(c_{2}\varepsilon)\right]\geqslant 0,$ (3.14)
where $K_{3}:=2^{\frac{n-1}{2}}K_{2}$. Rearranging (3.14), we conclude
$(C_{0}\varepsilon)^{-\frac{2}{n-1}}-K_{1}\left[\mathcal{H}\left(K_{3}T_{\varepsilon}^{-\frac{n-1}{2}}\varepsilon\right)-\mathcal{H}(c_{2}\varepsilon)\right]\geqslant 0$
$\qquad\Leftrightarrow\ K_{1}^{-1}(C_{0}\varepsilon)^{-\frac{2}{n-1}}+\mathcal{H}(c_{2}\varepsilon)\geqslant\mathcal{H}(K_{3}T_{\varepsilon}^{-\frac{n-1}{2}}\varepsilon)$
$\qquad\Leftrightarrow\ \mathcal{H}^{-1}\left(K_{1}^{-1}(C_{0}\varepsilon)^{-\frac{2}{n-1}}+\mathcal{H}(c_{2}\varepsilon)\right)\leqslant K_{3}T_{\varepsilon}^{-\frac{n-1}{2}}\varepsilon$
$\qquad\Leftrightarrow\ (K_{3}\varepsilon)^{\frac{2}{n-1}}\left[\mathcal{H}^{-1}\left(K_{1}^{-1}(C_{0}\varepsilon)^{-\frac{2}{n-1}}+\mathcal{H}(c_{2}\varepsilon)\right)\right]^{-\frac{2}{n-1}}\geqslant T_{\varepsilon},$
which is (2.7) with $c_{1}:=K_{1}^{-1}C_{0}^{-\frac{2}{n-1}}$ and $C:=K_{3}^{\frac{2}{n-1}}$. We stress that (2.7) is valid only for $\varepsilon\in(0,\varepsilon_{0}]$ for some $\varepsilon_{0}=\varepsilon_{0}(n,R,u_{1})>0$. Indeed, in (3.13) we have to assume $t\geqslant 2R$ and this condition is compatible with (2.7) only if we work with $\varepsilon\in(0,\varepsilon_{0}]$.
4 Upper bound estimates for the lifespan
In this section, we unravel the upper bound estimate (2.7) in the cases $(1\mathrm{MoC})$ and $(2\mathrm{MoC})$ from Example 2.1. We begin, by providing a rougher estimate than the one in (2.7) that we will use in some cases instead of (2.7) to better emphasize the dominant $\varepsilon$-dependent term in the upper bound for $T_{\varepsilon}$.
We remark that, by using de L’Hopital rule, $\mathcal{H}^{\prime}(\tau)=-\frac{\mu(\tau)}{\tau}$ and the continuity of $\mu(\tau)$ in $\tau=0$, we have
$\lim_{\varepsilon\to 0^{+}}\frac{\mathcal{H}(c_{2}\varepsilon)}{c_{1}\varepsilon^{-\frac{2}{n-1}}}=\lim_{\varepsilon\to 0^{+}}\frac{c_{2}\mathcal{H}^{\prime}(c_{2}\varepsilon)}{-\frac{2}{n-1}c_{1}\varepsilon^{-\frac{n+1}{n-1}}}=\lim_{\varepsilon\to 0^{+}}\frac{n-1}{2c_{1}}\dfrac{\dfrac{\mu(c_{2}\varepsilon)}{\varepsilon}}{\varepsilon^{-\frac{n+1}{n-1}}}=\lim_{\varepsilon\to 0^{+}}\frac{n-1}{2c_{1}}\mu(c_{2}\varepsilon)\,\varepsilon^{p_{\mathrm{Gla}}(n)-1}=0.$
Therefore, $\mathcal{H}(c_{2}\varepsilon)=o(\varepsilon^{-\frac{2}{n-1}})$ as $\varepsilon\to 0^{+}$, namely, in (2.7) in the argument of $\mathcal{H}^{-1}$ the term $\mathcal{H}(c_{2}\varepsilon)$ is dominated by the term $c_{1}\varepsilon^{-\frac{2}{n-1}}$. Hence, we have
$T_{\varepsilon}\leqslant C\varepsilon^{\frac{2}{n-1}}\left[\mathcal{H}^{-1}\left(2c_{1}\varepsilon^{-\frac{2}{n-1}}\right)\right]^{-\frac{2}{n-1}}$ (4.15)
for $\varepsilon\in(0,\varepsilon_{0}]$ with $\varepsilon_{0}>0$ small enough.
Finally, we point out that, in our proof of Theorem 2.1, the only properties of the nonlinearity $g(u_{t})=|u_{t}|^{p_{\mathrm{Gla}}(n)}\mu(|u_{t}|)$ that we actually employed are the convexity and the monotonicity of $g$, the fact that $g(u_{t})$ is a local nonlinearity (together with the property of finite speed of propagation for the wave equation) and the integral condition in (2.3), regardless of the fact that $\mu$ is or not a modulus of continuity. In other words, our proof is formally valid even in the cases $\mu(\tau)=\tau^{-\alpha}$ for $\alpha\in(0,\frac{2}{n-1})$ and $\mu(\tau)=1$ (clearly, these $\mu$ are not moduli of continuity), which correspond to the usual power nonlinearity of derivative-type in the sub-critical case $g(u_{t})=|u_{t}|^{p}$, with $1<p<p_{\mathrm{Gla}}(n)$ and in the critical case $g(u_{t})=|u_{t}|^{p_{\mathrm{Gla}}(n)}$, respectively. According to these considerations, we show that (2.7) recovers the known sharp upper bound estimates for the lifespan of a local solution to (1.1) both in the sub-critical and in the critical case.
Example (1MoC)
In this case $\mu(\tau)=\big{(}\log\frac{1}{\tau}\big{)}^{-\gamma}$ for any $\tau\in(0,\tau_{0}]$, where $\gamma\in(0,1]$. Hence,
$\mathcal{H}(\tau)=\begin{cases}\displaystyle{\frac{1}{1-\gamma}\left[\left(\log\frac{1}{\tau}\right)^{1-\gamma}-\left(\log\frac{1}{\tau_{0}}\right)^{1-\gamma}\,\right]}&\mbox{if}\ \gamma\in(0,1),\\[10.00002pt] \displaystyle{\log\log\frac{1}{\tau}-\log\log\frac{1}{\tau_{0}}}&\mbox{if}\ \gamma=1,\end{cases}$
and
$\mathcal{H}^{-1}(\omega)=\begin{cases}\displaystyle{\exp\left\{-\left[(1-\gamma)\omega+\left(\log\frac{1}{\tau_{0}}\right)^{1-\gamma}\,\right]^{\frac{1}{1-\gamma}}\right\}}&\mbox{if}\ \gamma\in(0,1),\\[13.00005pt] \displaystyle{\exp\left[-\left(\log\frac{1}{\tau_{0}}\right)\mathrm{e}^{\omega}\right]}&\mbox{if}\ \gamma=1,\end{cases}$
By using (4.15) for $\gamma\in(0,1)$ and (2.7) for $\gamma=1$ we obtain
$T_{\varepsilon}\leqslant\begin{cases}\displaystyle{C\varepsilon^{\frac{2}{n-1}}\exp\left(c\,\varepsilon^{-\frac{2}{(n-1)(1-\gamma)}}\right)}&\mbox{if}\ \gamma\in(0,1),\\[8.00003pt] \displaystyle{C\varepsilon^{\frac{2}{n-1}}\exp\left[c_{1}\left(\log\frac{1}{\varepsilon}\right)\exp\left(c_{2}\,\varepsilon^{-\frac{2}{n-1}}\right)\right]}&\mbox{if}\ \gamma=1.\end{cases}$
We underline that if we use (2.7) for $\gamma\in(0,1)$, then, as $\gamma\to 0^{+}$ we recover the upper bound estimate $T_{\varepsilon}\lesssim\exp\left(c\,\varepsilon^{-\frac{2}{n-1}}\right)$ for the power nonlinearity with critical Glassey exponent.
Example (2MoC)
In this case $\mu(\tau)=\prod_{j=1}^{k-1}\big{(}\log^{[j]}\frac{1}{\tau}\big{)}^{-1}\big{(}\log^{[k]}\frac{1}{\tau}\big{)}^{-\gamma}$ for any $\tau\in(0,\tau_{0}]$, where $k\geqslant 2$ is an integer and $\gamma\in(0,1]$. Thus,
$\mathcal{H}(\tau)=\begin{cases}\displaystyle{\frac{1}{1-\gamma}\left[\left(\log^{[k]}\frac{1}{\tau}\right)^{1-\gamma}-\left(\log^{[k]}\frac{1}{\tau_{0}}\right)^{1-\gamma}\,\right]}&\mbox{if}\ \gamma\in(0,1),\\[10.00002pt] \displaystyle{\log^{[k+1]}\frac{1}{\tau}-\log^{[k+1]}\frac{1}{\tau_{0}}}&\mbox{if}\ \gamma=1,\end{cases}$
and
$\mathcal{H}^{-1}(\omega)=\begin{cases}\displaystyle{\exp\left\{-\exp^{[k-1]}\left[(1-\gamma)\omega+\left(\log^{[k]}\frac{1}{\tau_{0}}\right)^{1-\gamma}\,\right]^{\frac{1}{1-\gamma}}\right\}}&\mbox{if}\ \gamma\in(0,1),\\[10.00002pt] \displaystyle{\exp\left[-\exp^{[k]}\left(\omega+\log^{[k+1]}\frac{1}{\tau_{0}}\right)\right]}&\mbox{if}\ \gamma=1.\end{cases}$
Employing (4.15) for $\gamma\in(0,1)$ and (2.7) for $\gamma=1$ we obtain
$T_{\varepsilon}\leqslant\begin{cases}\displaystyle{C\varepsilon^{\frac{2}{n-1}}\exp\left[c_{2}\exp^{[k-1]}\left(c_{1}\,\varepsilon^{-\frac{2}{(n-1)(1-\gamma)}}\right)\right]}&\mbox{if}\ \gamma\in(0,1),\\[10.00002pt] \displaystyle{C\varepsilon^{\frac{2}{n-1}}\exp\left\{c_{2}\exp^{[k-1]}\left[\left(\log^{[k]}\frac{1}{\varepsilon}\right)\exp\left(c_{1}\,\varepsilon^{-\frac{2}{n-1}}\right)\right]\right\}}&\mbox{if}\ \gamma=1.\end{cases}$
Sub-critical power nonlinearity
Let us consider formally $\mu(\tau)=\tau^{-\alpha}$ with $\alpha\in(0,\frac{2}{n-1})$. This choice corresponds to the nonlinearity $|u_{t}|^{p}$ with $1<p:=p_{\mathrm{Gla}}(n)-\alpha<p_{\mathrm{Gla}}(n)$. Then,
$\mathcal{H}(\tau)=\frac{1}{\alpha}\left(\tau^{-\alpha}-\tau_{0}^{-\alpha}\right)\ \ \mbox{and}\ \ \mathcal{H}^{-1}(\omega)=\left(\alpha\omega+\tau_{0}^{-\alpha}\right)^{-\frac{1}{\alpha}}.$
By (2.7), using that $\varepsilon^{-\frac{2}{n-1}}$ dominates $\varepsilon^{-\alpha}$, we obtain
$T_{\varepsilon}\lesssim\varepsilon^{\frac{2}{n-1}-\frac{4}{\alpha(n-1)^{2}}}=\varepsilon^{-(\frac{1}{p-1}-\frac{n-1}{2})^{-1}},$
which coincides with the sharp lifespan estimate (1.3) for the semilinear classical wave model (1.1) in the sub-critical Glassey case.
Critical power nonlinearity
Let us consider formally $\mu(\tau)=1$. In this case, the nonlinearity is exactly the power nonlinearity $|u_{t}|^{p_{\mathrm{Gla}}(n)}$. Thus,
$\mathcal{H}(\tau)=\log\tau_{0}-\log\tau\ \ \mbox{and}\ \ \mathcal{H}^{-1}(\omega)=\tau_{0}\,\mathrm{e}^{-\omega}.$
By (2.7) we find
$T_{\varepsilon}\lesssim\exp\left(c\,\varepsilon^{-\frac{2}{n-1}}\right),$
which coincides with the sharp lifespan estimate (1.3) for the semilinear classical wave model (1.1) in the critical Glassey case.
Acknowledgments
Wenhui Chen is supported in part by the National Natural Science Foundation of China (grant No. 12301270, grant No. 12171317), 2024 Basic and Applied Basic Research Topic–Young Doctor Set Sail Project (grant No. 2024A04J0016), Guangdong Basic and Applied Basic Research Foundation (grant No. 2023A1515012044). Alessandro Palmieri is a member of the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Instituto Nazionale di Alta Matematica (INdAM) and has been supported by INdAM - GNAMPA Project 2024 “Modelli locali e non-locali con perturbazioni non-lineari” CUP E53C23001670001 and by ERC Seeds UniBa Project “NWEinNES” CUP H93C23000730001. The authors thank Michael Reissig (TU Bergakademie Freiberg) for some suggestions in the preparation of the paper.